Air has a velocity of 1000 km/h at a pressure of 9.81 kN/m2 in vacuum

and a temperature of 47°C. Compute its stagnation properties and the local Mach number. Take

atmospheric pressure = 98.1 kN/m2, R = 287 J/kg K and γ = 1.4.

What would be the compressibility correction factor for a pitot-static tube to measure the

velocity at a Mach number of 0.8


1
Expert's answer
2021-09-16T14:19:04-0400

Velocity of air,

"V_0 = 1000 \;km/h = \frac{1000 \times 1000}{60 \times 60} = 277.78 \;m/s"

Temperature of air,

"T_0 = 47+273 = 320 \;K"

Atmospheric pressure,

"p_{atm} = 98.1 \;kN/m^2"

Pressure of air (static),

"p_0 = 98.1 -9.81=88.29 \;kN/m^2 \\

R=287 \;J/kg \;K \\

γ = 1.4"

Sonic velocity,

"C_0 = \sqrt{γ RT_0} = \sqrt{1.4 \times 287 \times 320} = 358.6 \;m/s"

Mach number,

"M_0 = \frac{V_0}{C_0} = \frac{277.78}{358.6} = 0.7746"

Stagnation pressure, ps :

The stagnation pressure is given by,

"p_s=p_0[1 +(\frac{γ-1}{2}M_0^2)]^{γ/(γ-1)} \\

p_s = 88.29[1 +(\frac{1.4-1}{2} \times 0.7746^2)]^{1.4/(1.4-1)} \\

= 88.29(1.12)^{3.5} = 131.27 \;kN/m^2"

Stagnation temperature, Ts :

"T_s =T_0[1 +(\frac{γ-1}{2})M_0^2] \\

T_s = 320[1 +(\frac{1.4-1}{2}) \times 0.7746^2] = 358.4 \;K \;or \;85.4 \;°C"

Stagnation density, ρs :

"ρ_s = \frac{p_s}{RT_s} = \frac{131.27 \times 10^3}{287 \times 358.4} = 1.276 \;kg/m^3"

Compressibility factor at M = 0.8 :

Compressibility factor "= 1 + \frac{M_0^2}{4} + \frac{2-γ}{24}M_0^4…"

"= 1 + \frac{0.8^2}{4} + \frac{2-1.4}{24} \times 0.8^4 = 1.1702"